How do I go about solving this? The answer is yes but I don't know why. Please help. Thank You very much:)
Determine whether the given point is in the solution set to the given system.
(-3, -1, 4)
2x-2y-3z=-16
4x-3y+3z=3
x+y-5z=-24

Answer by stanbon(75887)

Replace the x, the y, and the z with the values given to
see if you get three true statements:
---------------
Determine whether the given point is in the solution set to the given system.
(-3, -1, 4)
2x-2y-3z=-16
2*-3 -2*-1 -3*4 = -16
-6 +2 -12 = -16
-18+2 = -16
-16 = -16
So the 1st equation is true when x=-3,y=-1,z=4
-----------------------------------
Make the same substitution for the 2nd equation and then the 3rd:
4x-3y+3z=3
x+y-5z=-24
===========================
Cheers,
Stan H.